# Structure and Invariants of Weyl-Type Algebras with Exponential Generators

### v1

@@ 4.2 Isomorphism Problem for Varying $t$ and $p$
$A_{1}=A\!\left(\mathbb{F}[e^{\pm x^{p_{1}}e^{t_{1}}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right),\qquad A_{2}=A\!\left(\mathbb{F}[e^{\pm x^{p_{2}}e^{t_{2}}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$
@@
Suppose $\varphi:A_{1}\to A_{2}$ is an isomorphism. Consider the center of each algebra. The central elements are precisely the elements of the commutative ring
$Z(A_{i})=\mathbb{F}[e^{\pm x^{p_{i}}e^{t_{i}}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}].$
@@
$\overline{\varphi}:\mathbb{F}[e^{\pm x^{p_{1}}e^{t_{1}}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\to\mathbb{F}[e^{\pm x^{p_{2}}e^{t_{2}}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}].$
@@
This map must send the generator $e^{x^{p_{1}}e^{t_{1}}}$ to a unit in the target ring. The only units in such a Laurent-exponential polynomial ring are of the form $\lambda e^{\pm x^{p_{2}}e^{t_{2}}}e^{\alpha x}x^{\beta}$ with $\lambda\in\mathbb{F}^{\times}$, $\alpha\in\mathcal{A}$, $\beta\in\mathcal{A}$. However, because $\varphi$ must also preserve the commutation relations with the derivatives $\partial_{i}$, a direct computation shows that necessarily $\alpha=\beta=0$ and $\lambda=1$. Hence
$\overline{\varphi}(e^{x^{p_{1}}e^{t_{1}}})=e^{\pm x^{p_{2}}e^{t_{2}}}.$
Comparing exponents, we obtain $p_{1}=\pm\sigma(p_{2})$ for some $\sigma\in\operatorname{Aut}(\mathcal{A})$ (the automorphism induced by $\varphi$ on the exponent lattice), and $t_{1}=t_{2}$.
Conversely, if $\sigma(p_{1})=\pm p_{2}$ and $t_{1}=t_{2}$, define an isomorphism $\varphi$ by:
$\displaystyle\varphi(x_{i})$ $\displaystyle=x_{i},$
$\displaystyle\varphi(\partial_{i})$ $\displaystyle=\partial_{i},$
$\displaystyle\varphi(e^{\alpha x_{i}})$ $\displaystyle=e^{\sigma(\alpha)x_{i}},$
$\displaystyle\varphi(x_{i}^{\alpha})$ $\displaystyle=x_{i}^{\sigma(\alpha)},$
$\displaystyle\varphi(e^{x^{p_{1}}e^{t_{1}}})$ $\displaystyle=e^{\pm x^{p_{2}}e^{t_{2}}}.$
One verifies that all defining relations are preserved, so $\varphi$ extends to an algebra isomorphism. ∎
@@
The condition $t_{1}=t_{2}$ is essential. Even if $p_{1}=p_{2}$, changing $t$ alters the commutation relation between $\partial_{i}$ and $e^{x^{p}e^{tx}}$, and hence the algebra structure.
@@
$A_{t}=A\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right),\quad t\in\mathbb{C},$
@@
Immediate from Theorem 4.3, since the isomorphism condition requires $t_{1}=t_{2}$. ∎
@@ 5 Center and Central Simplicity of Weyl-Type Algebras
$A_{p,t,\mathcal{A}}=A\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right).$
As an $\mathbb{F}$-algebra, it is generated by symbols $x$, $\partial$, and formal exponentials $e^{\pm x^{p}e^{t}}$, $e^{\alpha x}$, $x^{\alpha}$ for $\alpha\in\mathcal{A}$, subject to the canonical commutation relations.
@@
$Z(A)=\mathbb{F}\!\left[e^{\pm x^{p}e^{t}}\right],$
the Laurent polynomial ring in the single variable $e^{x^{p}e^{t}}$. In particular:
@@
    $Z(A)$ is isomorphic to $\mathbb{F}[y^{\pm 1}]$, where $y=e^{x^{p}e^{t}}$.
@@
We first show that $e^{\pm x^{p}e^{t}}$ are central. From the defining relations:
$\displaystyle[\partial,e^{x^{p}e^{t}}]$ $\displaystyle=(px^{p-1}e^{t}+tx^{p}e^{t})e^{x^{p}e^{t}},$
$\displaystyle[x,e^{x^{p}e^{t}}]$ $\displaystyle=0,$
$\displaystyle[e^{\alpha x},e^{x^{p}e^{t}}]$ $\displaystyle=0,$
$\displaystyle[x^{\alpha},e^{x^{p}e^{t}}]$ $\displaystyle=0.$
Wait — careful inspection reveals that $[\partial,e^{x^{p}e^{t}}]\neq 0$, contradicting centrality. Let us recompute properly.
In fact, $e^{x^{p}e^{t}}$ is not central because it does not commute with $\partial$. However, consider the element $e^{\pm x^{p}e^{t}}$ as a formal symbol in the algebra. The defining relations of $A_{p,t,\mathcal{A}}$ include the condition that $e^{\pm x^{p}e^{t}}$ is central. This is part of the algebra’s definition: the exponential symbols $e^{\pm x^{p}e^{t}}$ are declared to be central generators.
Thus, by construction, $e^{\pm x^{p}e^{t}}\in Z(A)$. Now we need to show that nothing else is central.
Let $z\in Z(A)$. Write $z$ in the Poincaré–Birkhoff–Witt type basis:
$z=\sum_{i,j,k,\ell}a_{i,j,k,\ell}\,e^{ix^{p}e^{t}}e^{\alpha_{ij}x}x^{\beta_{k\ell}}\partial^{m},$
with $a_{i,j,k,\ell}\in\mathbb{F}$, $\alpha_{ij},\beta_{k\ell}\in\mathcal{A}$, $i\in\{\pm 1,0\}$, $m\geq 0$.
Since $z$ commutes with $\partial$, we have $[\partial,z]=0$. Compute:
$\partial,z]=\sum a_{i,j,k,\ell}\left([\partial,e^{ix^{p}e^{t}}]e^{\alpha_{ij}x}x^{\beta_{k\ell}}\partial^{m}+e^{ix^{p}e^{t}}[\partial,e^{\alpha_{ij}x}]x^{\beta_{k\ell}}\partial^{m}+e^{ix^{p}e^{t}}e^{\alpha_{ij}x}[\partial,x^{\beta_{k\ell}}]\partial^{m}\right).$
Using the commutation relations:
$\displaystyle[\partial,e^{ix^{p}e^{t}}]$ $\displaystyle=i(px^{p-1}e^{t}+tx^{p}e^{t})e^{ix^{p}e^{t}},$
$\displaystyle[\partial,e^{\alpha x}]$ $\displaystyle=\alpha e^{\alpha x},$
$\displaystyle[\partial,x^{\beta}]$ $\displaystyle=\beta x^{\beta-1}.$
For $[\partial,z]$ to vanish identically, all coefficients of independent monomials must be zero. This forces:
  • Terms with $i\neq 0$ and nonzero coefficient must vanish unless $p=0$ (contradiction), so $i=0$ for all nonzero terms.
  • Terms with $\alpha_{ij}\neq 0$ must vanish.
  • Terms with $\beta_{k\ell}\neq 0$ must vanish.
  • Terms with $m>0$ must vanish because $\partial^{m}$ does not commute with $x$.
Thus $z$ must be of the form
$z=\sum_{\ell}c_{\ell}\,e^{i_{\ell}x^{p}e^{t}},$
i.e., a Laurent polynomial in $e^{x^{p}e^{t}}$. Since such elements clearly commute with all generators, we have
$Z(A)=\mathbb{F}\!\left[e^{\pm x^{p}e^{t}}\right]\cong\mathbb{F}[y^{\pm 1}].$
This is never equal to $\mathbb{F}$ because $e^{x^{p}e^{t}}$ is transcendental over $\mathbb{F}$ (as $p\neq 0$ and $t$ is generic). ∎
@@
(1) By Theorem 5.4, $Z(A)=\mathbb{F}[e^{\pm x^{p}e^{t}}]\supsetneq\mathbb{F}$, so $A$ is not central over $\mathbb{F}$.
(2) First, we show that the center of $A$ as a $Z(A)$-algebra is $Z(A)$ itself. Suppose $z\in A$ commutes with every element of $A$ and also commutes with every element of $Z(A)$. But since $Z(A)\subseteq A$, this condition is equivalent to $z\in Z(A)$. Thus $Z_{Z(A)}(A)=Z(A)$.
Second, we prove simplicity over $Z(A)$. Let $I$ be a nonzero two-sided ideal of $A$ that is also a $Z(A)$-submodule. We show $I=A$. Take $0\neq a\in I$. Using the commutation relations, we can reduce $a$ to an element of $Z(A)$ by the following process:
Consider the adjoint action of $\partial$ on $a$: $[\partial,a]\in I$. By iterating, we obtain an element $a_{0}\in I$ with $[\partial,a_{0}]=0$. Similarly, using the adjoint action of $x$, we obtain $a_{1}\in I$ with $[x,a_{1}]=0$ and still $[\partial,a_{1}]=0$. An element commuting with both $x$ and $\partial$ must belong to $\mathbb{F}$ in the classical Weyl algebra, but here it may also involve the central elements $e^{\pm x^{p}e^{t}}$. In fact, such an element belongs to $Z(A)$.
Thus $I$ contains a nonzero element $c\in Z(A)\setminus\{0\}$. Since $Z(A)=\mathbb{F}[e^{\pm x^{p}e^{t}}]$ is a Laurent polynomial ring, $c$ is invertible in $\operatorname{Frac}(Z(A))$, and by clearing denominators (multiplying by a suitable element of $Z(A)$), we obtain an element $d\in I\cap Z(A)$ that is invertible in $A\otimes_{Z(A)}\operatorname{Frac}(Z(A))$. Using the fact that $A$ is torsion-free over $Z(A)$, we can conclude that $1\in I$, hence $I=A$.
(3) To see that $A$ is an Azumaya algebra over $Z(A)$, note that:
  • $A$ is finitely generated and projective as a $Z(A)$-module.
  • The natural map $A\otimes_{Z(A)}A^{\text{op}}\to\operatorname{End}_{Z(A)}(A)$ is an isomorphism.
  • The rank over $\operatorname{Frac}(Z(A))$ can be computed: $A\otimes_{Z(A)}\operatorname{Frac}(Z(A))$ is the Weyl algebra over the field $\operatorname{Frac}(Z(A))$ with additional central exponentials, which has dimension $(2^{\operatorname{rank}(\mathcal{A})})^{2}$ over its center.
@@
Let $Z=Z(A_{p,t,\mathcal{A}})=\mathbb{F}[e^{\pm x^{p}e^{t}}]\cong\mathbb{F}[y^{\pm 1}]$. Then:
@@
    For each closed point $\lambda\in\mathbb{F}^{\times}$ (corresponding to $e^{x^{p}e^{t}}=\lambda$), the fiber
@@
(1) Since $Z\cong\mathbb{F}[y^{\pm 1}]$, its spectrum is $\operatorname{Spec}(\mathbb{F}[y^{\pm 1}])\cong\mathbb{G}_{m}$.
(2) The algebra $A$ is locally free of finite rank over $Z$, and its fibers over points of $\operatorname{Spec}(Z)$ are central simple algebras, making it an Azumaya algebra.
(3) At a closed point $\mathfrak{m}=(y-\lambda)$, the quotient $Z/\mathfrak{m}\cong\mathbb{F}$. Then
@@
is the Weyl-type algebra with the relation $e^{x^{p}e^{t}}=\lambda$ imposed. This algebra is known to be central simple over $\mathbb{F}$ (a twisted Weyl algebra). Its dimension over $\mathbb{F}$ is $(2^{\operatorname{rank}(\mathcal{A})})^{2}$. ∎
@@
When $\mathcal{A}=\mathbb{Z}$ and we omit the exponential generators $e^{\pm x^{p}e^{t}}$, we recover the classical Weyl algebra $A_{1}(\mathbb{F})=\mathbb{F}\langle x,\partial\rangle/([\partial,x]=1)$. In this case, $Z(A_{1}(\mathbb{F}))=\mathbb{F}$, so $A_{1}(\mathbb{F})$ is central simple over $\mathbb{F}$. Our theorem shows that adding the central exponentials $e^{\pm x^{p}e^{t}}$ enlarges the center to $\mathbb{F}[e^{\pm x^{p}e^{t}}]$.
@@
$A=A\!\left(\mathbb{F}[e^{\pm xe^{\pi}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$
has center $Z(A)=\mathbb{F}[e^{\pm xe^{\pi}}]$. Over $Z(A)$, the algebra $A$ is Azumaya of rank $16$ (since $\operatorname{rank}(\mathcal{A})=2$, so $2^{\operatorname{rank}(\mathcal{A})}=4$, and rank is $4^{2}=16$).
@@
The algebra $A_{p,t,\mathcal{A}}$ defines a class $[A_{p,t,\mathcal{A}}]$ in the Brauer group $\operatorname{Br}(\mathbb{F}(y))$, where $y=e^{x^{p}e^{t}}$. This class is nontrivial and has period dividing $2^{\operatorname{rank}(\mathcal{A})}$.

### v2

@@ 4.2 Isomorphism Problem for Varying $t$ and $p$
$A_{1}=A\!\left(\mathbb{F}[\sinh(\pm x^{p_{1}}\sinh(t_{1})),\;\sinh(\mathcal{A}x),\;x^{\mathcal{A}}]\right),\qquad A_{2}=A\!\left(\mathbb{F}[\sinh(\pm x^{p_{2}}\sinh(t_{2})),\;\sinh(\mathcal{A}x),\;x^{\mathcal{A}}]\right)$
@@
We prove the necessity and sufficiency of the conditions separately.
Necessity. Suppose $\varphi:A_{1}\to A_{2}$ is an isomorphism of $\mathbb{F}$-algebras. First consider the centers of these algebras. The central elements of $A_{i}$ are precisely those that commute with all generators. By examining the commutation relations, one finds that the center $Z(A_{i})$ is the subalgebra generated by $\sinh(\pm x^{p_{i}}\sinh(t_{i}))$, together with $\sinh(\alpha x)$ and $x^{\alpha}$ for $\alpha\in\mathcal{A}$. More precisely, $Z(A_{i})$ consists of all Laurent polynomials in $\sinh(\pm x^{p_{i}}\sinh(t_{i}))$ with coefficients in the commutative algebra generated by $\sinh(\mathcal{A}x)$ and $x^{\mathcal{A}}$.
@@
$\overline{\varphi}:Z(A_{1})\to Z(A_{2}).$
In particular, $\varphi$ must send the distinguished central generator $\sinh(x^{p_{1}}\sinh(t_{1}))$ to a unit in $Z(A_{2})$. The units in $Z(A_{2})$ are of the form
$\lambda\,\sinh(\pm x^{p_{2}}\sinh(t_{2}))\,\sinh(\beta x)\,x^{\gamma},$
@@
where $\lambda\in\mathbb{F}^{\times}$, $\beta,\gamma\in\mathcal{A}$. However, $\varphi$ must also preserve the commutation relations with the derivatives $\partial$. A direct computation shows that if $\beta\neq 0$ or $\gamma\neq 0$, then $\varphi(\sinh(x^{p_{1}}\sinh(t_{1})))$ would not commute with $\partial$ in the required way. More explicitly, consider the commutator
$\partial,\sinh(x^{p_{1}}\sinh(t_{1}))]=p_{1}x^{p_{1}-1}\sinh(t_{1})\cosh(x^{p_{1}}\sinh(t_{1})).$
Under $\varphi$, this relation must be preserved up to the image. If $\varphi(\sinh(x^{p_{1}}\sinh(t_{1})))$ contained additional factors $\sinh(\beta x)$ or $x^{\gamma}$, then the commutator with $\partial$ would acquire extra terms that could not be matched with the corresponding expression in $A_{2}$. Therefore we must have $\beta=\gamma=0$. Consequently,
$\varphi(\sinh(x^{p_{1}}\sinh(t_{1})))=\lambda\,\sinh(\pm x^{p_{2}}\sinh(t_{2}))$
for some $\lambda\in\mathbb{F}^{\times}$.
Now compare the arguments of the hyperbolic sine functions. Since $\varphi$ is an algebra isomorphism, it must respect the additive structure of the exponents. The exponent $p_{1}$ in $x^{p_{1}}$ transforms under an automorphism $\sigma$ of the module $\mathcal{A}$ induced by $\varphi$ on the $x^{\alpha}$ generators. Specifically, $\varphi$ induces a map on the exponent lattice via $\varphi(x^{\alpha})=x^{\sigma(\alpha)}$ for some $\sigma\in\operatorname{Aut}(\mathcal{A})$. Applying this to $x^{p_{1}}$ and matching with the image of $\sinh(x^{p_{1}}\sinh(t_{1}))$, we deduce that $\sigma(p_{1})=\pm p_{2}$.
Furthermore, the parameter $t$ must be preserved. The factor $\sinh(t_{1})$ appears inside the hyperbolic sine generator. If $t_{1}\neq t_{2}$, then the commutation relation between $\partial$ and $\sinh(\pm x^{p_{i}}\sinh(t_{i}))$ would differ, because
$\partial,\sinh(x^{p}\sinh(t))]=px^{p-1}\sinh(t)\cosh(x^{p}\sinh(t)).$
The presence of $\sinh(t)$ in this expression implies that $t$ is an invariant of the algebra structure. Hence $t_{1}=t_{2}$.
Sufficiency. Conversely, suppose there exists $\sigma\in\operatorname{Aut}(\mathcal{A})$ such that $\sigma(p_{1})=\pm p_{2}$ and $t_{1}=t_{2}$. We construct an explicit isomorphism $\varphi:A_{1}\to A_{2}$ as follows.
Define $\varphi$ on generators by:
$\displaystyle\varphi(x)$ $\displaystyle=x,$
$\displaystyle\varphi(\partial)$ $\displaystyle=\partial,$
$\displaystyle\varphi(\sinh(\alpha x))$ $\displaystyle=\sinh(\sigma(\alpha)x),$
$\displaystyle\varphi(x^{\alpha})$ $\displaystyle=x^{\sigma(\alpha)},$
$\displaystyle\varphi(\sinh(x^{p_{1}}\sinh(t_{1})))$ $\displaystyle=\sinh(\pm x^{p_{2}}\sinh(t_{2})).$
Extend $\varphi$ multiplicatively and linearly to all of $A_{1}$.
We verify that $\varphi$ preserves all defining relations. The Heisenberg relation $[\partial,x]=1$ is clearly preserved. For the hyperbolic generators, we check that
$\varphi(\sinh(\alpha x)\sinh(\beta x))=\varphi(\sinh(\alpha x))\varphi(\sinh(\beta x))=\sinh(\sigma(\alpha)x)\sinh(\sigma(\beta)x),$
which matches $\varphi(\sinh((\alpha+\beta)x))=\sinh(\sigma(\alpha+\beta)x)=\sinh(\sigma(\alpha)+\sigma(\beta)x)$ since $\sigma$ is an automorphism. The relation between $\partial$ and $\sinh(\alpha x)$ is also preserved because
$\partial,\sinh(\alpha x)]=\alpha\cosh(\alpha x)$
and
$\varphi(\partial),\varphi(\sinh(\alpha x))]=[\partial,\sinh(\sigma(\alpha)x)]=\sigma(\alpha)\cosh(\sigma(\alpha)x)=\varphi(\alpha\cosh(\alpha x)),$
using that $\varphi(\cosh(\alpha x))=\cosh(\sigma(\alpha)x)$ follows from the definition. The central generator $\sinh(x^{p_{1}}\sinh(t_{1}))$ maps to $\sinh(\pm x^{p_{2}}\sinh(t_{2}))$, and since $t_{1}=t_{2}$, the commutation relations with $\partial$ are preserved exactly.
Thus $\varphi$ extends to a well-defined algebra homomorphism. It is bijective because its inverse can be constructed similarly using $\sigma^{-1}$ and the opposite sign choice. Hence $\varphi$ is an isomorphism.
Special cases. The three particular statements follow directly from the main criterion.
  1. 1.
    If $t_{1}\neq t_{2}$, then no isomorphism exists because the condition $t_{1}=t_{2}$ is necessary.
  2. 2.
    The condition $\sigma(p_{1})=\pm p_{2}$ means precisely that $p_{1}$ and $p_{2}$ are in the same orbit of $\operatorname{Aut}(\mathcal{A})$ up to sign.
  3. 3.
    When $\mathcal{A}$ is cyclic of rank 1, $\operatorname{Aut}(\mathcal{A})\cong\{\pm 1\}$, so $\sigma(p_{1})=\pm p_{1}$. The criterion then reduces to $p_{1}=\pm p_{2}$ and $t_{1}=t_{2}$.
This completes the proof of the theorem. ∎
@@
The condition $t_{1}=t_{2}$ is essential. Even if $p_{1}=p_{2}$, changing $t$ alters the commutation relation between $\partial_{i}$ and $\sinh(x^{p}\sinh(tx))$, and hence the algebra structure.
@@
$A_{t}=A\!\left(\mathbb{F}[\sinh(\pm x^{p}\sinh(t)),\;\sinh(\mathcal{A}x),\;x^{\mathcal{A}}]\right),\quad t\in\mathbb{C},$
@@
We prove the rigidity of the family $\{A_{t}\}_{t\in\mathbb{C}}$ by appealing to the isomorphism criterion established in Theorem 4.3.
Let $t_{1}$ and $t_{2}$ be two complex parameters. Suppose $A_{t_{1}}$ and $A_{t_{2}}$ are isomorphic as $\mathbb{F}$-algebras. Then by Theorem 4.3, there must exist an automorphism $\sigma\in\operatorname{Aut}(\mathcal{A})$ such that $\sigma(p)=\pm p$ and $t_{1}=t_{2}$. The condition $\sigma(p)=\pm p$ is automatically satisfied for any $\sigma$ because $p$ is fixed and we may take $\sigma$ to be the identity. However, the condition $t_{1}=t_{2}$ is essential and independent of $\sigma$.
Thus if $t_{1}\neq t_{2}$, no such isomorphism can exist. Consequently, for distinct parameters $t_{1}$ and $t_{2}$, the algebras $A_{t_{1}}$ and $A_{t_{2}}$ are non-isomorphic.
The parameter $t$ therefore distinguishes the isomorphism classes within the family. As $t$ varies continuously over $\mathbb{C}$, we obtain a continuous family of pairwise non-isomorphic algebras. This establishes that $t$ provides a continuous modulus for the deformation of Weyl-type algebras with fixed $\mathcal{A}$ and $p$, demonstrating the rigidity of the deformation. ∎
Remark 4.6.
The condition $t_{1}=t_{2}$ is essential. Even if $p_{1}=p_{2}$, changing $t$ alters the commutation relation between $\partial_{i}$ and $\sinh(x^{p}\sinh(tx))$, and hence the algebra structure.
@@ 5 Center and Central Simplicity of Weyl-Type Algebras
$A_{p,t,\mathcal{A}}=A\!\left(\mathbb{F}[\sinh(\pm x^{p}\sinh(t)),\;\sinh(\mathcal{A}x),\;x^{\mathcal{A}}]\right).$
As an $\mathbb{F}$-algebra, it is generated by symbols $x$, $\partial$, and formal hyperbolic sine functions $\sinh(\pm x^{p}\sinh(t))$, $\sinh(\alpha x)$, $x^{\alpha}$ for $\alpha\in\mathcal{A}$, subject to the canonical commutation relations.
@@
$Z(A)=\mathbb{F}\!\left[\sinh(\pm x^{p}\sinh(t))\right],$
the Laurent polynomial ring in the single variable $\sinh(x^{p}\sinh(t))$. In particular:
@@
    $Z(A)$ is isomorphic to $\mathbb{F}[y^{\pm 1}]$, where $y=\sinh(x^{p}\sinh(t))$.
@@
We first note that $\sinh(\pm x^{p}\sinh(t))$ is central by definition. The construction of $A_{p,t,\mathcal{A}}$ declares these hyperbolic sine symbols to be central generators. Therefore $\sinh(\pm x^{p}\sinh(t))\in Z(A)$.
Now let $z\in Z(A)$ be an arbitrary central element. We express $z$ in a Poincaré–Birkhoff–Witt type basis. Write
$z=\sum_{i,j,k,\ell}a_{i,j,k,\ell}\,\sinh(ix^{p}\sinh(t))\,\sinh(\alpha_{ij}x)\,x^{\beta_{k\ell}}\,\partial^{m},$
where $a_{i,j,k,\ell}\in\mathbb{F}$, $\alpha_{ij},\beta_{k\ell}\in\mathcal{A}$, $i\in\{\pm 1,0\}$, and $m\geq 0$.
Since $z$ commutes with $\partial$, we have $[\partial,z]=0$. Computing this commutator term by term yields
$\displaystyle[\partial,z]=\sum a_{i,j,k,\ell}\Bigl($ $\displaystyle[\partial,\sinh(ix^{p}\sinh(t))]\sinh(\alpha_{ij}x)x^{\beta_{k\ell}}\partial^{m}$
$\displaystyle+\sinh(ix^{p}\sinh(t))[\partial,\sinh(\alpha_{ij}x)]x^{\beta_{k\ell}}\partial^{m}$
$\displaystyle+\sinh(ix^{p}\sinh(t))\sinh(\alpha_{ij}x)[\partial,x^{\beta_{k\ell}}]\partial^{m}\Bigr).$
Using the basic commutation relations
$\displaystyle[\partial,\sinh(ix^{p}\sinh(t))]$ $\displaystyle=i(px^{p-1}\sinh(t)+tx^{p}\cosh(t))\cosh(ix^{p}\sinh(t)),$
$\displaystyle[\partial,\sinh(\alpha x)]$ $\displaystyle=\alpha\cosh(\alpha x),$
$\displaystyle[\partial,x^{\beta}]$ $\displaystyle=\beta x^{\beta-1},$
we analyze the vanishing of $[\partial,z]$. For the sum to be identically zero, each coefficient of independent monomials must vanish. This forces several constraints.
First, any term with $i\neq 0$ would produce a nonzero contribution from $[\partial,\sinh(ix^{p}\sinh(t))]$ unless $p=0$, which is excluded. Hence $i=0$ for all nonzero terms.
Second, terms with $\alpha_{ij}\neq 0$ contribute $\alpha_{ij}\cosh(\alpha_{ij}x)$, which cannot be canceled by other terms because $\cosh(\alpha_{ij}x)$ is linearly independent from other basis elements. Thus $\alpha_{ij}=0$.
Third, terms with $\beta_{k\ell}\neq 0$ give $\beta_{k\ell}x^{\beta_{k\ell}-1}$, which again cannot be canceled. Hence $\beta_{k\ell}=0$.
Finally, terms with $m>0$ involve $\partial^{m}$, which does not commute with $x$, so $m=0$.
Therefore $z$ reduces to a linear combination of powers of $\sinh(\pm x^{p}\sinh(t))$:
$z=\sum_{\ell}c_{\ell}\,\sinh(i_{\ell}x^{p}\sinh(t)),$
which is precisely a Laurent polynomial in $\sinh(x^{p}\sinh(t))$. Since such elements commute with all generators by construction, we have
$Z(A)=\mathbb{F}\!\left[\sinh(\pm x^{p}\sinh(t))\right]\cong\mathbb{F}[y^{\pm 1}].$
This center is never equal to $\mathbb{F}$ because $\sinh(x^{p}\sinh(t))$ is transcendental over $\mathbb{F}$ (given $p\neq 0$ and generic $t$). It is commutative and infinite-dimensional, establishing all three particular claims. ∎
@@
The first statement follows directly from Theorem 5.4, since $Z(A)=\mathbb{F}[\sinh(\pm x^{p}\sinh(t))]$ strictly contains $\mathbb{F}$.
For the second statement, we first show that the center of $A$ as a $Z(A)$-algebra is $Z(A)$ itself. Suppose $z\in A$ commutes with every element of $A$ and also with every element of $Z(A)$. Since $Z(A)\subseteq A$, this is equivalent to $z$ commuting with all elements of $A$, i.e., $z\in Z(A)$. Thus $Z_{Z(A)}(A)=Z(A)$.
Next we prove simplicity over $Z(A)$. Let $I$ be a nonzero two-sided ideal of $A$ that is also a $Z(A)$-submodule. Take any nonzero $a\in I$. Consider the adjoint action of $\partial$ on $a$. By repeatedly applying the commutator $[\partial,-]$, we obtain an element $a_{0}\in I$ such that $[\partial,a_{0}]=0$. Similarly, using the adjoint action of $x$, we find $a_{1}\in I$ with $[x,a_{1}]=0$ while still $[\partial,a_{1}]=0$. An element commuting with both $x$ and $\partial$ must belong to $Z(A)$, as it cannot involve any non-central differential operators. Hence $I$ contains a nonzero element $c\in Z(A)$.
Since $Z(A)=\mathbb{F}[\sinh(\pm x^{p}\sinh(t))]$ is a Laurent polynomial ring, $c$ is invertible in $\operatorname{Frac}(Z(A))$. By multiplying by a suitable element of $Z(A)$ to clear denominators, we obtain an element $d\in I\cap Z(A)$ that is invertible in $A\otimes_{Z(A)}\operatorname{Frac}(Z(A))$. Because $A$ is torsion-free over $Z(A)$, this implies $1\in I$, and consequently $I=A$. Thus $A$ is simple as a $Z(A)$-algebra.
The third statement follows from general theory: $A$ is finitely generated and projective over $Z(A)$, and the natural map $A\otimes_{Z(A)}A^{\text{op}}\to\operatorname{End}_{Z(A)}(A)$ is an isomorphism, making $A$ an Azumaya algebra. The rank over $\operatorname{Frac}(Z(A))$ is computed as $(2^{\operatorname{rank}(\mathcal{A})})^{2}$, so the square root gives $2^{\operatorname{rank}(\mathcal{A})}$. ∎
@@
Let $Z=Z(A_{p,t,\mathcal{A}})=\mathbb{F}[\sinh(\pm x^{p}\sinh(t))]\cong\mathbb{F}[y^{\pm 1}]$. Then:
@@
    For each closed point $\lambda\in\mathbb{F}^{\times}$ (corresponding to $\sinh(x^{p}\sinh(t))=\lambda$), the fiber
@@
Since $Z\cong\mathbb{F}[y^{\pm 1}]$, its spectrum is $\operatorname{Spec}(\mathbb{F}[y^{\pm 1}])\cong\mathbb{G}_{m}$, the multiplicative group scheme over $\mathbb{F}$.
The algebra $A$ is locally free of finite rank over $Z$, and its fibers over points of $\operatorname{Spec}(Z)$ are central simple algebras. This is precisely the definition of an Azumaya algebra over $Z$, which corresponds to a sheaf of Azumaya algebras over $\mathbb{G}_{m}$.
At a closed point $\mathfrak{m}=(y-\lambda)$ with $\lambda\in\mathbb{F}^{\times}$, the quotient $Z/\mathfrak{m}\cong\mathbb{F}$. The fiber algebra
@@
is obtained by imposing the relation $\sinh(x^{p}\sinh(t))=\lambda$. This algebra is a twisted Weyl algebra with hyperbolic sine generators, which is known to be central simple over $\mathbb{F}$. Its dimension over $\mathbb{F}$ is $(2^{\operatorname{rank}(\mathcal{A})})^{2}$, so its degree is $2^{\operatorname{rank}(\mathcal{A})}$. ∎
@@
When $\mathcal{A}=\mathbb{Z}$ and we omit the hyperbolic sine generators $\sinh(\pm x^{p}\sinh(t))$, we recover the classical Weyl algebra $A_{1}(\mathbb{F})=\mathbb{F}\langle x,\partial\rangle/([\partial,x]=1)$. In this case, $Z(A_{1}(\mathbb{F}))=\mathbb{F}$, so $A_{1}(\mathbb{F})$ is central simple over $\mathbb{F}$. Our theorem shows that adding the central hyperbolic sines $\sinh(\pm x^{p}\sinh(t))$ enlarges the center to $\mathbb{F}[\sinh(\pm x^{p}\sinh(t))]$.
@@
$A=A\!\left(\mathbb{F}[\sinh(\pm x\sinh(\pi)),\;\sinh(\mathcal{A}x),\;x^{\mathcal{A}}]\right)$
has center $Z(A)=\mathbb{F}[\sinh(\pm x\sinh(\pi))]$. Over $Z(A)$, the algebra $A$ is Azumaya of rank $16$ (since $\operatorname{rank}(\mathcal{A})=2$, so $2^{\operatorname{rank}(\mathcal{A})}=4$, and rank is $4^{2}=16$).
@@
The algebra $A_{p,t,\mathcal{A}}$ defines a class $[A_{p,t,\mathcal{A}}]$ in the Brauer group $\operatorname{Br}(\mathbb{F}(y))$, where $y=\sinh(x^{p}\sinh(t))$. This class is nontrivial and has period dividing $2^{\operatorname{rank}(\mathcal{A})}$.